OpenStudy (anonymous):

Ben invested an amount of money at 5.2% per annum compound interest.? Calculate the number of years after which the value of his investment had doubled,

OpenStudy (anonymous):

@Nnesha '

OpenStudy (anonymous):

@babyblue123

OpenStudy (anonymous):

@Jensen

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

\[A = P(1 + \frac{R}{100})^T\] R = 5.2 A = 2P [Because the amount doubles] Find T. :D

OpenStudy (anonymous):

can u solve it please @AkashdeepDeb i stopped at 2=(1.052)^t but i don't know what to do next??

OpenStudy (akashdeepdeb):

Do you have a calculator?

OpenStudy (akashdeepdeb):

And do you know what are logarithms by any chance? ;-)

OpenStudy (anonymous):

yes i do but i didn't take logarithms so if there is any other method??